Calculate the amount of heat observed from the heat reservoir by the engine in a cycle, if the engine does work of 200J to the surroundings and release heat of 800J to the surroundings in the cycle. Also calculate the efficiency of the engine that works in that cycle.
1
Expert's answer
2012-04-05T09:46:13-0400
The amount of heat in a cycle consists of heat released to the surroundings and of work to the surroundings that transforms into heat. So, the amount of heat is 200J+800J = 1000J. The efficiency of the engine is the ratio of useful work to released heat: μ = 200J/800J = 0.25 (or 25%).